If x^2 - 5x +4 = 0 , then x =
Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

If xsquared - 5X +4 = 0 , then x =
----
x^2-5x+4 = 0
Factor:
(x-4)(x-1) = 0
x = 4 or x = 1
====================
Cheers,
Stan H.

Answer by ewatrrr(24785)

 
Hi,
x^2 - 5x +4 = 0
factoring
(x-4)(x-1) = 0
(x-4)= 0, x = 4
OR(x-1) = 0, x = 1
Note Graph of y = x^2 - 5x +4